OpenStudy (anonymous):

Ugh sorry guys.. stuck on this one now 6i/-1-3i can someone show me steps? Thought I had these down but I suppose I do not

OpenStudy (anonymous):

6i ___ -1-3i

OpenStudy (anonymous):

\[\frac{6i}{-1-3i}=\frac{-6i}{1+3i}\times\frac{1-3i}{1-3i}=\frac{-6i+18(-1)}{1-9(-1)}=\frac{18-6i}{10}=\frac{9}{5}-\frac{3i}{5}\]

OpenStudy (anonymous):

ooppss typo ...

OpenStudy (anonymous):

I dropped the "-" sign from "18" in the end ... The answer should be: -9/5 - 3i/5

OpenStudy (anonymous):

Ah, I see it

OpenStudy (anonymous):

Thank you so very much for helping with this, I appreciate it.. for some reason I'm having a lot of trouble with these >.<

OpenStudy (anonymous):

The mistake was intentional - just checking if you were alert ...

OpenStudy (anonymous):

lol Appreciate you keeping me on my toes ;)